OpenStudy (anonymous):

2y^2-5y-12=0 how would you solve by completing the square?

OpenStudy (wolf1728):

Move the "non X" term to the right: 2y² -5y = 12 Divide the equation by the coefficient of X² which in this case is 2 y² -2.5y = 6 (There's more)

OpenStudy (wolf1728):

take the coefficient of y which is -2.5 divide it by 2 -1.25 square that number -1.25² = 1.5625 then add it to both sides of the equation. y² -2.5y + 1.5625= 7.5625 Take the square root of both sides of the equation: y -1.25 = sqrt(7.5625) y = 1.25 + 2.75 y = 4 sqrt(7.5625) also equals -2.75 y -1.25 = -2.75 y = -1.5 Source for completing the square: http://1728.org/quadr2.htm